OpenStudy (anonymous):

Which of the following is a factor of x^2 – 8x – 48? x – 8 x + 4 x – 6 None of the above

OpenStudy (anonymous):

x^2 – 8x – 48 = x^2 -12x + 4x - 48 = x(x - 12) + 4(x - 12) = (x + 4)(x - 12) therefore x+4 is a factor of x^2 – 8x – 48